# EmotiCrafter: Text-to-Emotional-Image Generation based on Valence-Arousal Model

### v2

@@ Abstract
00footnotetext: Shengqi Dang and Yi He contributed equally to this work.00footnotetext: Nan Cao is the corresponding author.00footnotetext: Shengqi Dang, Yi He, Long Ling, Ziqing Qian, and Nan Cao are with the Intelligent Big Data Visualization Lab, Tongji University. Shengqi Dang and Nan Cao are also with the Shanghai Innovation Institution. Email: {dangsq123, heyi_11}@tongji.edu.cn, lucyling0224@gmail.com, 2411920@tongji.edu.cn, nan.cao@gmail.com.00footnotetext: Nanxuan Zhao is with Adobe Research. Email: nanxuanzhao@g- mail.com.00footnotetext: Code: https://github.com/idvxlab/EmotiCrafter
@@ 1 Introduction
Emotions are fundamental to human experiences and play a critical role in shaping how people perceive and interact with the world. Research has shown that emotions affect memory [19, 15, 25, 45] and comprehension [16, 40, 38], which are crucial for effective communication. As a result, content creators increasingly recognize the importance of incorporating emotions to enhance audience engagement.
While research on visual emotion analysis is extensive [49, 13, 58], there is limited work on generating emotionally rich image content. Some early studies explored emotional content generation techniques within specific domains such as facial expressions [2, 46] or landscapes [24]. EmoGen [50] generates images based on a given emotion tag (e.g., happy or sad), which pioneered the domain-free emotional image content generation (EICG) task. However, it has two critical limitations: (1) the image is generated from an emotion tag instead of a text prompt, making the generated content difficult to control; (2) while the discrete emotion tags used in EmoGen are easy to understand, psychologists have not achieved a consensus on the emotion categories [60]. The limited scope of discrete emotion tags falls short of capturing nuanced emotions.
To address the above issues, we propose the continuous emotional image content generation (C-EICG) task and present EmotiCrafter, the first C-EIGC model that generates emotional images using free-text prompts and continuous Valence-Arousal (V-A) values defined in the V-A model (a well-known psychological continuous emotion model)  [33]. V-A model represents emotions in a two-dimensional Cartesian space (Figure 2), where Valence quantifies pleasantness (negative to positive) and Arousal measures intensity (calm to excited). We utilize a [-3, 3] range for Valence and Arousal [10, 21]. The continuous V-A space enables smooth transitions and nuanced emotional shifts beyond the capability of discrete labels (e.g., shifts from “bored” to “tired” of a character’s state in a video). Specific emotional values in this space have been investigated in prior work [34], and such fine-grained modeling is particularly beneficial for human–computer interaction [2]. Leveraging this model, EmotiCrafter captures subtle affective variations via precise (V, A) positioning. The contributions of this paper are as follows:
@@
  • We propose a novel task, continuous emotional image content generation (C-EICG), and develop the first dedicated model for this purpose. Our model introduces an emotion-embedding network that integrates continuous Valence-Arousal (V-A) values with text prompts. These fused features are then injected into Stable Diffusion XL [28] using cross-attention mechanisms, enabling precise control over both content and emotional expressions in the generated images.
@@ 2.1 Visual Emotion Analysis
Visual emotion analysis refers to the computational recognition and interpretation of human emotion [57] in visual media, such as images or videos. It has been a prominent research area, with most efforts focusing on the classification of discrete emotions [52, 48, 42, 4, 47]. However, discrete emotion categories limit the ability to capture the nuanced emotions [7], leading to an increased focus on continuous emotion analysis in images [59].
Much of the continuous emotion analysis remains centered on facial expression analysis [37, 22, 54, 9, 11]. While effective, facial analysis can overlook crucial contextual information that influences emotional interpretation. This drives studies to focus on emotions in objects or individuals within their environment, rather than just faces[21, 12, 20, 13].
For example, Kosti et al. [12] combined person-specific characteristics with the context of the scene to predict continuous emotional dimensions such as arousal, valence, and dominance. Kragel et al. [13] proposed EmoNet to extract visual features such as facial expressions, body posture, and scene elements from images to predict valence and arousal values. Recently, Mertens et al.[21] tested multiple backbones (such as ResNet, CLIP, DINO, etc.) for prediction of valence and arousal, demonstrating strong performance across various architectures.
@@ 2.2 Image Emotion Transfer
Image Emotion Transfer (IET) focuses on editing the content of the images to evoke different emotions [17, 61, 27, 43]. For instance, Peng et al. [27] achieved emotion transfer by adjusting color tones and texture-related features. Zhu et al. [61] introduced a method that separates high-level emotion-relevant features (e.g., object shapes and scene layout) from low-level emotion-relevant features (e.g., brightness). By applying GANs, they transferred emotions between images while preserving their original structure. Building on these methods, Weng et al. [43] proposed the Affective Image Filter, which uses a multi-modal transformer to process both text and image inputs. With the emergence of text-to-image models, IET has expanded into new applications based on instructive commands. For example, EmoEdit [51] used GPT-4V to build emotion factor trees that map abstract emotions to specific visual elements and employed the InstructPix2Pix model to apply emotion-driven content and color adjustments to images.
@@ 2.3 Conditional Image Generation
Conditional image generation aims to create images that align with specific input conditions, such as text [31, 28, 30, 3], reference images [53], subjects [6, 32], and depth maps [55, 23]. To enhance quality, researchers have developed specialized approaches, such as Diffusion Transformers (DiT)[26], which use transformer-based diffusion for denoising, and Visual Autoregressive Modeling (VAR)[36], which encodes images into discrete tokens for autoregressive prediction across scales. However, these state-of-the-art methods rely on discrete labels, limiting flexibility and control. In contrast, our approach offers greater generalization by leveraging Stable Diffusion XL [28] to generate emotional images from free-text prompts.
Despite these progresses, incorporating emotion as a condition for image generation remains underexplored. EmoGen [50] pioneered the generation of emotional image content (EIGC) by mapping emotional features to semantic features to generate emotional images. However, EmoGen struggles to understand natural language, limiting its capacity to effectively control specific content. Its reliance on discrete emotions also limits its practical applicability.
@@ 3.2 Emotion-Embedding Network
The emotion-embedding network $\mathcal{M}$ generates the emotional prompt feature by integrating a pair of V-A values with a neutral prompt feature (Fig.3(b.1)). First, a V-A encoder converts the V-A values into feature vectors. Then, an emotion injection transformer—modified from GPT-2[29]—fuses these vectors with the neutral prompt feature, preserving the original textual context while infusing emotional content.
@@
where $h_{0}$ represents the initial hidden state; $\text{P}_{\text{in}}(\cdot)$ is a linear projection layer, and PE denotes positional embedding [39].
@@
where $h_{i}$ is the output of the $i$-th EIB()($\cdot$)( ⋅ ). As shown in Figure 4, each EIB enhances the transformer block through a cross-attention mechanism:
@@
[…]
@@
V-A Density Weighting. To mitigate the effects of the imbalanced distribution of the training samples, we weigh the loss inversely proportional to the density of training samples in the V-A space. The density is estimated using Kernel Density Estimation (KDE) [5] with a Gaussian kernel, denoted as $d(v,a)$:
@@ 3.3 Dataset and Training
To train the emotion-embedding network, we constructed a dataset of paired neutral and emotional prompts with corresponding Valence-Arousal (V-A) values (Figure 3(a)). These pairs were automatically generated using GPT-4 based on 39,843 images with human-annotated V-A values from publicly available datasets, including OASIS [14], EMOTIC [12], and FindingEmo [21]. Specifically, GPT-4 was used to generate neutral prompts with objective image descriptions and emotional prompts that emphasize affective attributes such as color, lighting, and texture, which influence emotional perception. To ensure data reliability, all LLM-generated prompts underwent crowd-sourced verification, with disagreements resolved through a voting mechanism among annotators.
The proposed emotion embedding network is trained on two NVIDIA A800 GPUs using the aforementioned dataset. We employ the AdamW [18] optimizer with a weight decay of 1e-5 and a learning rate of 1e-3. The training process spans 200 epochs with a batch size of 768, completing in approximately 7 to 8 hours.
@@ 4.2 Comparisons
As a result, four different baselines were built: (1) Cross Attention: inject the emotion features ($e_{v},e_{a}$) into the UNet in SDXL via its cross-attention mechanism based on IP-Adapter [53]; (2) Time Embedding: directly add emotion features ($e_{v},e_{a}$) to the time embedding of the UNet in SDXL. (3) Textual Inversion: use the text inversion technique [6] to embed emotion features ($e_{v},e_{a}$) into prompt templates with predefined emotion placeholders. (4) GPT-4+SDXL (GPT-SD): use GPT-4 [1] to rewrite the input text according to (v,a) values to generate an emotional SDXL prompt for image generation.
@@
Quantitative Comparison. We compare our method against several baselines using the following metrics: (1) V/A-Error evaluates the absolute error between the predicted V/A of the generated images and the input V/A. (2) CLIPScore [8] assesses the similarity between the input text and the generated images. (3) CLIP-IQA [41] leverages a pre-trained CLIP model to evaluate image quality without requiring reference images. (4) LPIPS-Continuous utilizes the Learned Perceptual Image Patch Similarity [56] to measure the continuity of the change in image as V/A changes.
@@
Cross Attention 1.923$\pm$1.153 2.080$\pm$1.438 26.266$\pm$2.381 0.949$\pm$0.046
Time Embedding 1.941$\pm$1.168 2.031$\pm$1.348 26.566$\pm$2.125 0.786$\pm$0.164
Textual Inversion 1.958$\pm$1.188 1.923$\pm$1.170 22.346$\pm$3.594 0.370$\pm$0.111
GPT-4+SDXL 1.860$\pm$1.090 1.517$\pm$1.060 25.907$\pm$1.949 0.906$\pm$0.066
Ours 1.828$\pm$1.085 1.510$\pm$1.074 23.067$\pm$2.655 0.881$\pm$0.099
@@ 4.3 User Study
Analysis & Results. We conducted a Shapiro-Wilk test [35] to assess normality and applied the Wilcoxon Signed Rank test [44] to evaluate statistical significance. Our results indicate that our method outperformed the baseline across all metrics (Table 3), with statistically significant improvements in V/A Ranking, V/A Error, Consistency, and Smoothness (Figure 8).
@@ 5 Conclusion and Limitations
Although our method achieved promising results, it still has some limitations need to be addressed in the future. First, controlling image generation based on arousal remains more challenging than controlling based on valence. This is consistent with prior research in visual emotion analysis, which has found that arousal is harder to predict due to lower inter-annotator agreement [21]. Second, our approach frequently generates images featuring human activities even when such activities are not mentioned in the prompts. This likely stems from the limited representation of non-human scenes in our training data and could be mitigated by incorporating a more diverse range of non-human scenarios. Third, our method occasionally modifies users’ input prompts to better align with the specified emotional prompts, resulting in a slight semantic shift that affects the generated images. We believe this issue could be addressed by adding a semantic preservation term to the loss function.

### v3

@@ Abstract
00footnotetext: Shengqi Dang and Yi He contributed equally to this work.00footnotetext: Nan Cao is the corresponding author.00footnotetext: Shengqi Dang, Yi He, Long Ling, Ziqing Qian, and Nan Cao are with the Intelligent Big Data Visualization Lab, Tongji University. Shengqi Dang and Nan Cao are also with the Shanghai Innovation Institute. Email: {dangsq123, heyi_11}@tongji.edu.cn, lucyling0224@gmail.com, 2411920@tongji.edu.cn, nan.cao@gmail.com.00footnotetext: Nanxuan Zhao is with Adobe Research. Email: nanxuanzhao@g- mail.com.00footnotetext: Code: https://github.com/idvxlab/EmotiCrafter
@@ 1 Introduction
Emotions are fundamental to human experiences and play a critical role in shaping how people perceive and interact with the world. Research has shown that emotions affect memory [megalakaki2019effects, lee1999effects, pawlowska2011influence, xie2017negative] and comprehension [li2020can, vega1996representation, tyng2017influences], which are crucial for effective communication. As a result, content creators increasingly recognize the importance of incorporating emotions to enhance audience engagement.
While research on visual emotion analysis is extensive [yang2017joint, kragel2019emotion, zhao2014exploring], there is limited work on generating emotionally rich image content. Some early studies explored emotional content generation techniques within specific domains such as facial expressions [azari2024emostyle, xu2023high] or landscapes [park2020emotional]. EmoGen [yang2024emogen] generates images based on a given emotion tag (e.g., happy or sad), which pioneered the domain-free emotional image content generation (EICG) task. However, it has two critical limitations: (1) the image is generated from an emotion tag instead of a text prompt, making the generated content difficult to control; (2) while the discrete emotion tags used in EmoGen are easy to understand, psychologists have not achieved a consensus on the emotion categories [9472932]. The limited scope of discrete emotion tags falls short of capturing nuanced emotions.
To address the above issues, we propose the continuous emotional image content generation (C-EICG) task and present EmotiCrafter, the first C-EIGC model that generates emotional images using free-text prompts and continuous Valence-Arousal (V-A) values defined in the V-A model (a well-known psychological continuous emotion model)  [russell1980circumplex]. V-A model represents emotions in a two-dimensional Cartesian space (Figure 2), where Valence quantifies pleasantness (negative to positive) and Arousal measures intensity (calm to excited). We utilize a [-3, 3] range for Valence and Arousal [khota2022modelling, mertens2024findingemo]. The continuous V-A space enables smooth transitions and nuanced emotional shifts beyond the capability of discrete labels (e.g., shifts from “bored” to “tired” of a character’s state in a video). Specific emotional values in this space have been investigated in prior work [evidence], and such fine-grained modeling is particularly beneficial for human–computer interaction [azari2024emostyle]. Leveraging this model, EmotiCrafter captures subtle affective variations via precise (V, A) positioning. The contributions of this paper are as follows:
@@
  • We propose a novel task, continuous emotional image content generation (C-EICG), and develop the first dedicated model for this purpose. Our model introduces an emotion-embedding network that integrates continuous Valence-Arousal (V-A) values with text prompts. These fused features are then injected into Stable Diffusion XL [sdxl] using cross-attention mechanisms, enabling precise control over both content and emotional expressions in the generated images.
@@ 2.1 Visual Emotion Analysis
Visual emotion analysis refers to the computational recognition and interpretation of human emotion [zhang2024refashioning] in visual media, such as images or videos. It has been a prominent research area, with most efforts focusing on the classification of discrete emotions [9246699, 10.1007/978-3-031-19836-6_9, wang2022systematic, borth2013large, xu2022mdan]. However, discrete emotion categories limit the ability to capture the nuanced emotions [5771357], leading to an increased focus on continuous emotion analysis in images [10.1145/2733373.2806354].
Much of the continuous emotion analysis remains centered on facial expression analysis [toisoul2021estimation, 8013713, zhang2014representation, 9881519, Kollias_2022_CVPR]. While effective, facial analysis can overlook crucial contextual information that influences emotional interpretation. This drives studies to focus on emotions in objects or individuals within their environment, rather than just faces[mertens2024findingemo, Kosti_2017_CVPR, meng2022multi, kragel2019emotion]. For example, Kosti et al. [Kosti_2017_CVPR] combined person-specific characteristics with the context of the scene to predict continuous emotional dimensions such as arousal, valence, and dominance. Kragel et al. [kragel2019emotion] proposed EmoNet to extract visual features such as facial expressions, body posture, and scene elements from images to predict valence and arousal values. Recently, Mertens et al.[mertens2024findingemo] tested multiple backbones (such as ResNet, CLIP, DINO, etc.) for prediction of valence and arousal, demonstrating strong performance across various architectures.
@@ 2.2 Image Emotion Transfer
Image Emotion Transfer (IET) focuses on editing the content of the images to evoke different emotions [LIU201816, zhu2023emotional, peng2015mixed, weng2023affective]. For instance, Peng et al. [peng2015mixed] achieved emotion transfer by adjusting color tones and texture-related features. Zhu et al. [zhu2023emotional] introduced a method that separates high-level emotion-relevant features (e.g., object shapes and scene layout) from low-level emotion-relevant features (e.g., brightness). By applying GANs, they transferred emotions between images while preserving their original structure. Building on these methods, Weng et al. [weng2023affective] proposed the Affective Image Filter, which uses a multi-modal transformer to process both text and image inputs. With the emergence of text-to-image models, IET has expanded into new applications based on instructive commands. For example, EmoEdit [yang2024emoedit] used GPT-4V to build emotion factor trees that map abstract emotions to specific visual elements and employed the InstructPix2Pix model to apply emotion-driven content and color adjustments to images.
@@ 2.3 Conditional Image Generation
Conditional image generation aims to create images that align with specific input conditions, such as text [sd, sdxl, dalle2, dalle3], reference images [ip-adaptor], subjects [texutual-inverison, dreambooth], and depth maps [controlnet, mou2023t2i]. To enhance quality, researchers have developed specialized approaches, such as Diffusion Transformers (DiT)[dit], which use transformer-based diffusion for denoising, and Visual Autoregressive Modeling (VAR)[var], which encodes images into discrete tokens for autoregressive prediction across scales. However, these state-of-the-art methods rely on discrete labels, limiting flexibility and control. In contrast, our approach offers greater generalization by leveraging Stable Diffusion XL [sdxl] to generate emotional images from free-text prompts.
Despite these progresses, incorporating emotion as a condition for image generation remains underexplored. EmoGen [yang2024emogen] pioneered the generation of emotional image content (EIGC) by mapping emotional features to semantic features to generate emotional images. However, EmoGen struggles to understand natural language, limiting its capacity to effectively control specific content. Its reliance on discrete emotions also limits its practical applicability.
@@ 3.2 Emotion-Embedding Network
The emotion-embedding network $\mathcal{M}$ generates the emotional prompt feature by integrating a pair of V-A values with a neutral prompt feature (Fig.3(b.1)). First, a V-A encoder converts the V-A values into feature vectors. Then, an emotion injection transformer—modified from GPT-2[radford2019language]—fuses these vectors with the neutral prompt feature, preserving the original textual context while infusing emotional content.
@@
where $h_{0}$ represents the initial hidden state; $\text{P}_{\text{in}}(\cdot)$ is a linear projection layer, and PE denotes positional embedding [vaswani2017attention].
@@
where $h_{i}$ is the output of the $i$-th EIB()($\cdot$). As shown in Figure 4, each EIB enhances the transformer block through a cross-attention mechanism:
@@
where $h^{\prime},h^{(v)},h^{(v,a)}$ are the intermediate hidden variables; $\text{LN}(\cdot)$ is the LayerNorm; $\text{self-attn}(\cdot)$ denotes self-attention, employed to capture context dependencies; $\text{cross-attn}(\cdot)$ is cross-attention for injecting $e_{v}$ and $e_{a}$; $\text{fnn}(\cdot)$ is a feed-forward network that adapts the complexity of the emotional embedding process. We also remove the causal mask typically used for next-token prediction from the original transformer model to fit our task.
@@
V-A Density Weighting. To mitigate the effects of the imbalanced distribution of the training samples, we weigh the loss inversely proportional to the density of training samples in the V-A space. The density is estimated using Kernel Density Estimation (KDE) [chen2017tutorial] with a Gaussian kernel, denoted as $d(v,a)$:
@@ 3.3 Dataset and Training
To train the emotion-embedding network, we constructed a dataset of paired neutral and emotional prompts with corresponding Valence-Arousal (V-A) values (Figure 3(a)). These pairs were automatically generated using GPT-4 based on 39,843 images with human-annotated V-A values from publicly available datasets, including OASIS [OASIS], EMOTIC [Kosti_2017_CVPR], and FindingEmo [mertens2024findingemo]. Specifically, GPT-4 was used to generate neutral prompts with objective image descriptions and emotional prompts that emphasize affective attributes such as color, lighting, and texture, which influence emotional perception. To ensure data reliability, all LLM-generated prompts underwent crowd-sourced verification, with disagreements resolved through a voting mechanism among annotators.
The proposed emotion embedding network is trained on two NVIDIA A800 GPUs using the aforementioned dataset. We employ the AdamW [loshchilov2017decoupled] optimizer with a weight decay of 1e-5 and a learning rate of 1e-3. The training process spans 200 epochs with a batch size of 768, completing in approximately 7 to 8 hours.
@@ 4.2 Comparisons
As a result, four different baselines were built: (1) Cross Attention: inject the emotion features ($e_{v},e_{a}$) into the UNet in SDXL via its cross-attention mechanism based on IP-Adapter [ip-adaptor]; (2) Time Embedding: directly add emotion features ($e_{v},e_{a}$) to the time embedding of the UNet in SDXL. (3) Textual Inversion: use the text inversion technique [texutual-inverison] to embed emotion features ($e_{v},e_{a}$) into prompt templates with predefined emotion placeholders. (4) GPT-4+SDXL (GPT-SD): use GPT-4 [gpt4] to rewrite the input text according to (v,a) values to generate an emotional SDXL prompt for image generation.
@@
Quantitative Comparison. We compare our method against several baselines using the following metrics: (1) V/A-Error evaluates the absolute error between the predicted V/A of the generated images and the input V/A. (2) CLIPScore [hessel2021clipscore] assesses the similarity between the input text and the generated images. (3) CLIP-IQA [clipiqa] leverages a pre-trained CLIP model to evaluate image quality without requiring reference images. (4) LPIPS-Continuous utilizes the Learned Perceptual Image Patch Similarity [zhang2018unreasonable] to measure the continuity of the change in image as V/A changes.
@@
[…]
@@ 4.3 User Study
Analysis & Results. We conducted a Shapiro-Wilk test [shaphiro1965analysis] to assess normality and applied the Wilcoxon Signed Rank test [wilcoxon1992individual] to evaluate statistical significance. Our results indicate that our method outperformed the baseline across all metrics (Table 3), with statistically significant improvements in V/A Ranking, V/A Error, Consistency, and Smoothness (Figure 8).
@@ 5 Conclusion and Limitations
Although our method achieved promising results, it still has some limitations need to be addressed in the future. First, controlling image generation based on arousal remains more challenging than controlling based on valence. This is consistent with prior research in visual emotion analysis, which has found that arousal is harder to predict due to lower inter-annotator agreement [mertens2024findingemo]. Second, our approach frequently generates images featuring human activities even when such activities are not mentioned in the prompts. This likely stems from the limited representation of non-human scenes in our training data and could be mitigated by incorporating a more diverse range of non-human scenarios. Third, our method occasionally modifies users’ input prompts to better align with the specified emotional prompts, resulting in a slight semantic shift that affects the generated images. We believe this issue could be addressed by adding a semantic preservation term to the loss function.